# Multi-Modal Discussion Transformer: Integrating Text, Images and Graph Transformers to Detect Hate Speech on Social Media

### 08 Jan 2024 v3

@@ Abstract
We present the Multi-Modal Discussion Transformer (mDT), a novel methodfor detecting hate speech in online social networks such as Reddit discussions. In contrast to traditional comment-only methods, our approach to labelling a comment as hate speech involves a holistic analysis of text and images grounded in the discussion context. This is done by leveraging graph transformers to capture the contextual relationships in the discussion surrounding a comment and grounding the interwoven fusion layers that combine text and image embeddings instead of processing modalities separately. To evaluate our work, we present a new dataset, HatefulDiscussions, comprising complete multi-modal discussions from multiple online communities on Reddit. We compare the performance of our model to baselines that only process individual comments and conduct extensive ablation studies.
@@ Results
We conduct a 7-fold stratified cross-validation with a fixed seed (1) and report the average performance for each model. We report overall accuracy (Acc.) and class-weighted Precision (Pre.), Recall (Rec.) and F1 to account for label imbalance. In all results, * denotes statistical significance using Student’s t-test with p-value $<0.05$. Our hyperparameter search space can be seen in Table 2. All experiments were run using 2xA100-40GB GPUs, a 12-core Intel CPU, 80GB of RAM, and Linux.
@@
Table 3: Performance of mDT against Text-Only Methods
@@
mDT (ours) 0.880* 0.880* 0.880* 0.877*
@@
Table 6: Effect of Fusion Layers
Total Fusion Layers Acc. Pre. Rec. F1
@@
Usage of Images Acc. Pre. Rec. F1
With Images 0.880* 0.880* 0.880* 0.877*
Without Images 0.832 0.835 0.822 0.828
@@
We next perform a qualitative comparison of the text-only BERT model and the proposed mDT architecture. We find that the text-only BERT model misclassifies 385/2717 test instances. Upon passing those test instances through mDT, we found that it corrected BERT’s labels in 161/385 instances. We further note that BERT and mDT predictions disagree on 264 test instances, which mDT is correct on 161 (61%). Figure 3 shows a fine-grained distribution of misclassified test examples by class. Using mDT results in an overall decrease in misclassifications (385 $\rightarrow$ 327), with a major reduction in false positives (fewer misclassifications for the ‘Not Hateful’ class). However, we notice that BERT and mDT both struggle to detect the presence of hate speech in derogatory slur (DEG) and identity-directed (IdentityDirectedAbuse) comments.

### 22 Feb 2024 v4

@@ Abstract
We present the Multi-Modal Discussion Transformer (mDT), a novel method for detecting hate speech on online social networks such as Reddit discussions. In contrast to traditional comment-only methods, our approach to labelling a comment as hate speech involves a holistic analysis of text and images grounded in the discussion context. This is done by leveraging graph transformers to capture the contextual relationships in the discussion surrounding a comment and grounding the interwoven fusion layers that combine text and image embeddings instead of processing modalities separately. To evaluate our work, we present a new dataset, HatefulDiscussions, comprising complete multi-modal discussions from multiple online communities on Reddit. We compare the performance of our model to baselines that only process individual comments and conduct extensive ablation studies.
@@ Results
We conduct a 7-fold stratified cross-validation with a fixed seed (1) and report the average performance for each model. We report overall accuracy (Acc.) and class-weighted Precision (Pre.), Recall (Rec.) and F1 to account for label imbalance. Underlined values denote statistical significance using Student’s t-test with p-value $<0.05$ in all results. Our hyperparameter search space can be seen in Table 2. All experiments were run using 2xA100-40GB GPUs, a 12-core Intel CPU, 80GB of RAM, and Linux.
@@
mDT (ours) 0.880 0.880 0.880 0.877
Table 3: Performance of mDT against Text-Only Methods
@@
Fusion Layers Acc. Pre. Rec. F1
@@
Usage of Images Acc. Pre. Rec. F1
With Images 0.880 0.880 0.880 0.877
Without Images 0.832 0.835 0.822 0.828
@@
We next perform a qualitative comparison of the text-only BERT model and the proposed mDT architecture. We find that the text-only BERT model misclassifies 385/2717 test instances. Upon passing those test instances through mDT, we found that it corrected BERT’s labels in 161/385 cases. We further note that BERT and mDT predictions disagree on 264 test instances, which mDT is correct on 161 (61%). Figure 3 shows a fine-grained distribution of misclassified test examples by class. Using mDT results in an overall decrease in misclassifications (385 $\rightarrow$ 327), with a significant reduction in false positives (fewer misclassifications for the ‘Not Hateful’ class). However, we notice that BERT and mDT both struggle to detect the presence of hate speech in derogatory slur (DEG) and identity-directed (IdentityDirectedAbuse) comments.